User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Katie

Attended: Jamie, Lucia, Patrick, Lucas, Simin, Jesse, Ross, Tiege, Atifur

Apologies: Mas

Absent: Prad

General Issues

  • Yg wind sensor?
    • Something in the wiring, disconnected at the moment (under investigation). Keep an eye on the weather radar. Cal Yg if you see wind picking up. Monica?
  • Ke mk5 problems?
    • Crashing and stalling. Maybe going to change it. Maybe swap to mk52. Under investigation.
  • AUM exp went well, have fringes.
  • Masers for ke and yg not working.
    • ke visible on one network but not the other…yg not sure. Under investigation.
  • Audio on ops8 - drops in and out.
  • Half of baby monitor missing.
  • Ho 26 L band receiver being repaired. Cold head being assessed in Mlb.
  • Ho 26 maybe one of the breaks has a problem?
  • hobart 12, for flogit, the proc files weren't getting sent. The proc file is essentially empty, its a 'dummy'.
  • occasionally an error with comments, and thats why flogit won't work sometimes.

Handover Issues

  • R4853 :
    • ke: ES Oven V Katherine maser fluctuate from blue colour (0.9) to red colour (25.8). Brett came to auscope room and had a look at this problem. We sent an email the history file of maser KE to Anton and Jamie (mas)
    • Open issues with the maser being investigated. Noone can go do some repairs inside the maser - another reason why it is important to keep an eye on the maser.
/home/www/auscope/opswiki/data/pages/operations/agenda2018_07_26.txt · Last modified: 2018/08/09 03:21 by Simin Salarpour